# Zero-shot cross-lingual transfer in instruction tuning of large language models

### v1

@@ Abstract
Instruction tuning (IT) is widely used to teach pretrained large language models (LLMs) to follow arbitrary instructions, but is under-studied in multilingual settings. In this work, we conduct a systematic study of zero-shot cross-lingual transfer in IT, when an LLM is instruction-tuned on English-only data and then tested on user prompts in other languages. We investigate the influence of model configuration choices and devise a multi-facet evaluation strategy for multilingual instruction following. We find that cross-lingual transfer does happen successfully in IT even if all stages of model training are English-centric, but only if multiliguality is taken into account in hyperparameter tuning and with large enough IT data. English-trained LLMs are capable of generating correct-language, comprehensive and helpful responses in the other languages, but suffer from low factuality and may occasionally have fluency errors.
@@ 1 Introduction
Most of the IT research has focused on English, leaving multilingual instruction following a rather understudied area. Several recent works aim to extend instruction tuning beyond English by creating target language IT datasets via automatic translation of English instructions Cab (2023); Zic (2023), distilling the outputs of powerful models such as GPT-4 (Wei et al., 2023), or crowdsourcing (Köpf et al., 2023; Singh et al., 2024). However, all of these strategies incur high costs or effort and require repeating the data creation process for each language of interest (target language).
Figure 1: Zero-shot cross-lingual transfer setting in instruction tuning: an LLM is instruction-tuned on English-only data and then tested on user prompts in other languages. Our study focuses on model configuration choices and analyzing various aspects of generated outputs.
In this work, we take a close look at zero-shot cross-lingual transfer in instruction tuning, when the LLM is tuned solely on English instruction data and then prompted to follow instructions in target languages without any additional target-language adaptation. Such an approach has the clear advantages of low cost and easy applicability to various target languages but is often considered just as a simple baseline, without detailed analysis. We aim to deeper understand (1) which factors influence the successful cross-lingual knowledge transfer as well as (2) what are the capabilities and limits of the zero-shot approach.
To answer the first question, we study a range of model configurations and investigate the influence of the base model (multilingual / English-centric, model size), IT data size, adaptation strategy and hyperparameters. To answer the second question we devise a multi-facet, fine-grained evaluation strategy of multilingual instruction following. The most common evaluation strategy consists of scoring the helpfulness of model responses on some publicly available set of diverse instructions, e.g. AlpacaFarm Dubois et al. (2023), with powerful model, e.g. GPT-3.5. The shortcomings of such high-level evaluation are lack of control over the distribution of tasks in the evaluation set, unclear interpretation of helpfulness scores (what aspects of the answers are wrong: fluency, content, relevance to the task etc), and uncertain reliability of LLMs as evaluators, especially in languages other than English. This motivates us to devise a more careful evaluation pipeline, which includes a diverse and balanced set of instructions and a variety of metrics (surface metrics, GPT-3.5 evaluation, and human evaluation) for various aspects of generated texts. Our key findings include:
@@ 2 Related work
Most of the works in multilingual IT aim to extend the IT dataset with non-English data (Köpf et al., 2023; Wei et al., 2023; Cab, 2023; Zic, 2023; Shaham et al., 2024; Kew et al., 2023; Chen et al., 2023; Ranaldi et al., 2023), or decompose non-English instructions by pivoting through English translations (Zhang et al., 2023b; Etxaniz et al., 2023). Chen et al. (2023); Kew et al. (2023); Shaham et al. (2024) advocate for the sufficiency of a "pinch" of multilinguality in IT, represented by a small amount of updates on multilingual IT data, small amount of multilingual IT data mixed with English data, or having only 2–3 languages in the IT data. We focus on English-only IT, trying to better assess capabilities and limits of such settings.
The concurrent work of Shaham et al. (2024) does demonstrate the proof-of-the-concept results on zero-shot cross-lingual transfer in IT, but attributes it to the multilinguality of PaLM-2 pretraining data. We show cross-lingual transfer in IT for English-centric models and conduct a more deep and systematic investigation of this effect.
@@
3 Methodology
The main goal of our study is to understand the capabilities and limitations of zero-shot cross-lingual transfer in instruction tuning of LLMs, and which factors influence them. We consider cases when the model is pretrained on English-centric and multilingual data. In the former case, the multilingual instruction-following capabilities of the model arise solely from the small amount of occasional multilingual data which is always present in English-centric pretraining corpora crawled from the Internet Blevins and Zettlemoyer (2022). We study the effect of various model choices such as the base model size, the size of the English instruction data, adaptation strategy (full or parameter-efficient finetuning), and adaptation hyperparameters.
Evaluation metrics. Our evaluation strategy includes three components: (1) easy-to-compute surface metrics, (2) manual inspection of predictions, and (3) evaluation with GPT-3.5.
@@ 3 Methodology
The surface metrics include language correctness (how often the language of the response matches the language of the task), spellcheck correctness (which portion of words in the responses pass spell checking), and relevance to the task (how often responses are relevant to their tasks, evaluated using LLama-2-chat-7B). These metrics serve to identify if a model passes a minimal bar on the quality of multilingual answers and help to select hyperparameters and filter out non-effective model configurations without GPT-based evaluation.
@@
After that, the main evaluation is conducted using both manual predictions inspection (on a subset of the evaluation set) and GPT-3.5 (on the full evaluation set). To control qualitative aspects of generated texts, we judge them with 6 criteria: helpfulness (how helpful in general is the response for the user), language correctness (does the language of the response match the language of the task), fluency, factual accuracy, logical coherence and harmlessness. Five of these criteria (except language correctness) were introduced in (Zhang et al., 2023a) and in our preliminary study we found that they reflect well the weaknesses of model responses. We also use the same scale from 0 to 2.
Controlling tasks distribution. We identify a diverse set of 25 "tasks" present in AlpacaFarm (e.g. write an email, give advice, rewrite text etc.) and select a subset of 113 instructions from AlpacaFarm that include a balanced number of instructions per "task". Thus obtained set of 113 instructions is used in GPT-3.5-based evaluation, and a stratified subset of 30 instructions is used in human evaluation. Controlling the task distribution ensures that none of the tasks dominates the evaluation set, leading to more reliable conclusions, and allows us to break down the performance results by tasks.
Control of task complexity. To deeper analyze the effect of task complexity, we introduce a set of task modifiers which add details to the task, such as generate a short or detailed response, answer in a specified language or style, format the answer in a specified way, or answer two questions one after another. Modifiers are manually translated into target languages and added to instructions one-by-one. We evaluate overall helpfulness of the produced responses (taking into account all given instructions) and modifier fullfilment: whether responses follow additional instructions given in the modifier.
4 Experimental setup
Base LLMs. In our work we consider (1) LLaMA-2 at 7B and 13B sizes, (2) TowerBase-7B, built on top of LLaMa-2-7B, further trained on balanced data covering 10 languages. This allows us to assess an importance of multilinguality at pretraining.
IT strategy. We consider two most popular supervised finetuning techniques: full finetuning and LoRA finetuning. We train all listed model configurations with three learning rates (LRs) and choose the best LR based on surface metrics.
Composite tasks. For each modifier we select a subset of 15-100 appropriate input instructions.
For more experimental details, see Appendix B.
Figure 3: Results of human evaluation (left) and evaluation with GPT-3.5 (right). All scores from 0 to 2. Base models: LLaMA-2-7B/13B (English-centric) or Tower-7B (10 languages). Datasets: Dolly (15k) or LIMA (1k). Data strategies: En (English-only data) or DT (multilingual data obtained using data translation). Adaptation strategy: FT (full finetuning) or LoRA (low-rank adaptation).
5 Experimental results and discussion
Preliminary study based on surface metrics.
Figure 2 (left) presents results for a wide range of model configurations, each trained with three LRs and evaluated using the surface metrics. Surface metrics help to choose optimal LRs and identify configurations which do not pass the minimal bar for cross-lingual transfer with any of the LRs.
@@ 5 Experimental results and discussion
Careful hyperparameter tuning and in particular LR selection is essential for achieving multilingual instruction following capabilities. All the model configurations, except training on the small LIMA data, achieve high values for all metrics in all languages with LR of 1e-5 (1e-4 for LoRa adaptation). The lower LR of 1e-6 leads to lower relevance scores in some languages, due to model under-training. On the other side, the higher LR of 1e-4 leads to overfitting to the training language(s), pronounced by lower language correctness scores and lower spellcheck correctness scores. The lower spellcheck correctness is caused by code-switching.
Even though training on small instruction data was shown to be enough for English (Zhou et al., 2023), it substantially reduces the cross-lingual capabilities of the final model compared to training on the larger data. Training on the small LIMA data leads to severe overfitting to English with all LRs, pronounced by low language correctness scores and low spellcheck correctness. In particular, this model produces a lot of incoherent texts and for morphology-rich Russian, generates a lot of made-up words formed from existing word roots, even with a smaller LR of 1e-5. Such made-up words can be present in other models as well but much less frequently.
Main evaluation.
Figure 3 shows the results of human (left) and GPT-3.5-based (right) evaluation, for English and average over Fr, Pt, and Ru. Per-language results are presented in App. Figure 4. We observe generally consistent trends between GPT-3.5 and human evaluation.
Instruction-tuned model is able to successfully transfer learned knowledge to other languages, but with helpfulness to some extent lower than in English. The main criteria, Helpfulness, achieves best score of 1.35 in non-English settings on average (out of 2, human evaluation) for the English-centric model, LLaMA-2-13B/Dolly-En/FT. It is slightly higher compared to LLaMA-2-13B/Dolly-DT/FT integrating multilinguality in the IT dataset, but lower compared to human-evaluated Helpfulness in English (1.77). Same trends hold for GPT-3.5 evaluation.
Factuality suffers most in cross-lingual transfer. The factual accuracy score is lower for non-English than for English in all cases, and is among the lowest criteria scores for all models, e.g. 1.46 for LLaMA-2-13B/Dolly-En/FT. Factual accuracy of baselines trained with multilingual data is not higher than of the English-centric model, posing a challenge for future works at improving truthfulness in the multilingual setting.
Fluency, logicality and correct language scores occasionally suffer in zero-shot cross-lingual transfer. Human-evaluated scores for correct language, fluency and logical coherence in non-English are around 1.8 out of 2 for the default model and slightly higher, around 1.9, for the multilingually pretrained Tower-7B/Dolly-En/FT. We highlight that the problem of generation in the wrong language appears rarely in cross-lingual setting (after careful LR tuning), opposite to the conclusions of prior work Chen et al. (2023).
Ablations (smaller IT dataset, small base LLM, LoRA adaptation) reduce helpfulness in non-English. Aligned with surface metrics, the model tuned on (English) LIMA is characterized by very low scores for all criteria, in non-English evaluation [^1] Absolute helpfulness score assigned by GPT-3.5, 1.52, is substantially higher than the one assigned in human evaluation, 0.94, because LIMA-based model produces much longer outputs than Dolly-based model and GPT-3.5 is known to be biased towards long verbose responses.. Scores for English are close to other models, which aligns with the initial findings of (Zhou et al., 2023). Using LoRA instead of full finetuning and decreasing model size reduce most of the scores.
Task modifier
Mod. fulfill. Helpfulness
en
ru
en
ru
Answer briefly in just a few sentences.
80%
90%
1.70
1.60
Give a detailed answer.
65%
75%
1.60
1.55
List N options (N random from 2 to 10)
66%
83%
1.66
1.66
Answer in X language. (X: Fr, Pt, De)
47%
79%
1.37
1.47
Use markdown formatting in the answer.
92%
100%
1.85
1.28
Format your answer as an html page.
57%
14%
1.35
1.00
Begin each point with the sign –>
7%
14%
0.92
0.85
Capitalize each first letter in the answer.
7%
7%
1.00
0.64
Write in a scientific style.
92%
92%
1.64
1.57
The answer should use simple words.
78%
78%
1.57
1.28
Two-hop instruction, e.g. explain how to serve a dish after telling how to cook it.
93%
86%
1.80
1.60
Average
62%
65%
1.49
1.31
Table 1: Performance with various task modifiers. Modifier fulfilness measures the percentage of inputs for which the modifier was fulfilled. Helpfulness also takes conditions specified in modifiers into account.
Helpfulness in non-English reduces in some language-related tasks, tasks involving calculation or US-centric factual knowledge. Figure 2 (right) breaks down human-evaluated helpfulness by task category. We find that English-centric model struggles in other languages with some of language-based tasks such as rewriting given sentences, suggesting words that rhyme with the given one or follow a given pattern. At the same time, models do succeed on easier language-related tasks such as generate synonyms or words beginning with a given letter. Models also make calculation errors more often in non-English than in English. The low helpfulness for the "sport game" category is connected to the low factuality in non-English: this category asks to explain rules of games popular in the USA and they are explained well in English and often hallucinated in other languages.
English-centric models are capable of following composite instructions in non-English languages in  65% of cases. Results in Table 1 show that the majority of task modifiers are fulfilled in around 80% of cases with helpfulness score being similar to the value observed in the main evaluation. Interestingly, instruction to generate response in another language, is fulfilled substantially more often when it is written in non English. An example of the instruction that often fails in non-English is to format the answer as an html page.

### v2

@@ Abstract
Instruction tuning (IT) is widely used to teach pretrained large language models (LLMs) to follow arbitrary instructions, but is under-studied in multilingual settings. In this work, we conduct a systematic study of zero-shot cross-lingual transfer in IT, when an LLM is instruction-tuned on English-only data and then tested on user prompts in other languages. We advocate for the importance of evaluating various aspects of model responses in multilingual instruction following and investigate the influence of different model configuration choices. We find that cross-lingual transfer does happen successfully in IT even if all stages of model training are English-centric, but only if multiliguality is taken into account in hyperparameter tuning and with large enough IT data. English-trained LLMs are capable of generating correct-language, comprehensive and helpful responses in the other languages, but suffer from low factuality and may occasionally have fluency errors.
@@ 1 Introduction
Most of the IT research has focused on English, leaving multilingual instruction following a rather understudied area. Several recent works aim to extend instruction tuning beyond English by creating target language IT datasets via automatic translation of English instructions Cab (2023); Zic (2023), distillation of outputs of powerful models such as GPT-4 (Wei et al., 2023; Li et al., 2023), or crowdsourcing (Köpf et al., 2023; Singh et al., 2024). However, all of these strategies incur high costs or effort and require repeating the data creation process for each language of interest (target language).
Figure 1: Zero-shot cross-lingual transfer in instruction tuning: an LLM is instruction-tuned on English-only data and then tested on user prompts in other languages. Our study focuses on analyzing various aspects of generated outputs and model configuration choices.
In this work, we take a close look at zero-shot cross-lingual transfer in instruction tuning, when the LLM is tuned solely on English instruction data and then prompted to follow instructions in target languages without any additional target-language adaptation. Such an approach has the clear advantages of low cost and easy applicability to various target languages but is often considered just as a simple baseline, without detailed analysis. We aim to deeper understand (RQ1) what are the capabilities and limits of the zero-shot approach as well as (RQ2) which factors influence the successful cross-lingual knowledge transfer.
The most common strategy for evaluating instruction following capabilities consists of scoring the helpfulness of model responses on some publicly available set of diverse instructions, e.g. AlpacaFarm Dubois et al. (2023), with a powerful model, e.g. GPT-3.5. We argue that such high-level evaluation is insufficient and not informative enough for a multi-facet task of open-ended generation, especially in the multilingual scenario. We advocate for using a more careful evaluation pipeline, including the evaluation of various aspects of model responses (fluency, content, relevance to the task etc.), controlling the distribution and the complexity of the tasks in the evaluation set, and using both automatic and human evaluation. This allows us to characterize the weak and strong sides of multilingual responses generated by the model tuned on English-only data (RQ1) and to better understand the influence of factors such as the base model (multilingual / English-centric, model size), IT data size, adaptation strategy and hyperparameters (RQ2). Our key findings include:
@@ 2 Related work
Most of the works in multilingual IT aim to extend the IT dataset with non-English data (Köpf et al., 2023; Singh et al., 2024; Li et al., 2023; Wei et al., 2023), or decompose non-English instructions by pivoting through English translations (Zhang et al., 2023b; Etxaniz et al., 2023). Chen et al. (2024); Kew et al. (2023); Shaham et al. (2024) advocate for the sufficiency of a "pinch" of multilinguality in IT, represented by a small amount of updates on multilingual IT data, small amount of multilingual IT data mixed with English data, or having only 2–3 languages in the IT data. We focus on English-only IT, trying to better assess capabilities and limits of such settings.
The concurrent work of Shaham et al. (2024) does demonstrate the proof-of-the-concept results on zero-shot cross-lingual transfer in IT, but attributes it to the multilinguality of PaLM-2 pretraining data. We show that cross-lingual transfer in IT works well even for English-centric models and conduct a more deep and systematic investigation of this effect.
@@
3 Our evaluation methodology
To better understand the strong and weak sides of multilingual responses generated by the model tuned on English-only data, we devise a multi-facet evaluation strategy which includes evaluation of various aspects of generated responses, controlling task distribution and complexity, and using both model-based and human evaluation.
Evaluation criteria. We conduct main evaluation using both manual predictions inspection (on a subset of the evaluation set) and GPT-3.5 evaluation (on the full evaluation set). To control qualitative aspects of generated texts, we judge them with 6 criteria: helpfulness (how helpful in general is the response for the user), language correctness (does the language of the response match the language of the task), fluency, factual accuracy, logical coherence and harmlessness. Five of these criteria (except language correctness) were introduced in (Zhang et al., 2023a) and in our preliminary study we found that they reflect well the weaknesses of model responses. We also use the same scale from 0 to 2 for each criteria.
@@ 3 Our evaluation methodology
We also introduce lightweight surface metrics: language correctness (how often the language of the response matches the language of the task), spellcheck correctness (which portion of words in the responses pass spell checking), and relevance to the task (how often responses are relevant to their tasks, evaluated using LLama-2-chat-7B). These metrics serve to identify if a model passes a minimal bar on the quality of multilingual answers and help to select hyperparameters and filter out non-effective model configurations without GPT-based evaluation.
Refer to captionRefer to caption
Figure 2: Results of human evaluation (left) and evaluation with GPT-3.5 (right). All scores from 0 to 2, heatmap colors visualize written scores. Base models: LLaMA-2-7B/13B (English-centric) or Tower-7B (10 languages). Datasets: Dolly (15k) or LIMA (1k). Instruction tuning data strategies: En (English-only data) or DT (multilingual IT data obtained using data translation). Adaptation strategy: FT (full finetuning) or LoRA (low-rank adaptation).
Control of the task distribution. We identify a diverse set of 25 "tasks" present in AlpacaFarm (e.g. write an email, give advice, rewrite text etc.) and select a subset of 113 instructions from AlpacaFarm that include a balanced number of instructions per "task". Thus obtained set of 113 instructions is used in GPT-3.5-based evaluation, and a stratified subset of 30 instructions is used in human evaluation. Controlling the task distribution ensures that none of the tasks dominates the evaluation set, leading to more reliable conclusions, and allows us to break down the performance results by tasks.
Control of the task complexity. To deeper analyze the effect of task complexity, we introduce a set of task modifiers which add details to the task, such as generate a short or detailed response, answer in a specified language or style, format the answer in a specified way, or answer two questions one after another. Modifiers are manually translated into target languages and added to instructions one-by-one. For each modifier we select a subset of 15-100 appropriate input instructions. We evaluate overall helpfulness of the produced responses (taking into account all given instructions) and modifier fulfillment: whether responses follow additional instructions given in the modifier.
4 Experimental setup
We study the effect of various choices such as the base model, the size of the English instruction data, adaptation strategy (full or parameter-efficient finetuning), and adaptation hyperparameters.
Base LLMs. In our work we consider (1) LLaMA-2 Touvron et al. (2023) at 7B and 13B sizes, (2) TowerBase-7B Alves et al. (2024), built on top of LLaMa-2-7B, further trained on balanced data covering 10 languages. In the former case, the multilingual instruction-following capabilities of the model arise solely from the small amount of occasional multilingual data which is always present in English-centric pretraining corpora crawled from the Internet Blevins and Zettlemoyer (2022). The latter case allows us to assess an importance of multilinguality at pretraining.
IT strategy. We consider two most popular supervised finetuning techniques: full finetuning (FT) and LoRA finetuning.
We select LLaMA-2-13B/Dolly-En/FT as an anchor model configuration and apply changes to it one-by-one, i.e. changing the base model, IT data, or the adaptation method. We train all model configurations with three learning rates (LRs) and choose the best LR based on surface metrics. For more experimental details, see Appendix B.
5 Experimental results and discussion
5.1 Main evaluation
Figure 2 shows the results of human (left) and GPT-3.5-based (right) evaluation, for English and average over Fr, Pt, and Ru. Per-language results are presented in App. Figure 4. Agreement between automatic and human evaluation is visualized in App. Figure 5. Though we observe generally consistent trends between GPT-3.5 and human evaluation in average scores, they can disagree in evaluating individual samples, especially for the scores of helpfulness, factual accuracy, and fluency. Agreement for non-English is lower than for English.
RQ1. We first analyze various aspects of predictions for our anchor English-centric and English-tuned model, LLaMA-2-13B/Dolly-En/FT.
Instruction-tuned model is able to successfully transfer learned knowledge to other languages, but with helpfulness to some extent lower than in English. The main score, overall Helpfulness, for our anchor English-centric model, LLaMA-2-13B/Dolly-En/FT, achieves 1.77 / 1.35 in English / non-English settings correspondingly (out of 2, human evaluation). As we discuss below, one of the main factors contributing to this difference is reduced factuality in non-English. Another factor is that responses in non-English sometimes contain obvious advice, e.g. "to install a window blind, follow the instructions provided with it" (translated from Russian).
Factuality is the weak side of predictions in non-English. The factual accuracy score is substantially lower in non-English than in English, e.g. 1.46 vs 1.80 in human evaluation. This poses a challenge for future works at improving truthfulness in the multilingual setting.
English-tuned model may occasionally (but rarely) produce output in the wrong language, code-switching, or make a fluency error. Scores for correct language, fluency and logical coherence are between 1.8 and 1.9 for the anchor model LLaMA-2-13B/Dolly-En/FT in non-English settings. This holds for both automatic and human evaluation, except GPT-3.5 evaluation of fluency, demonstrating the need for the better automatic evaluation of this criteria. We highlight that the problem of generation in the wrong language appears rarely in cross-lingual setting (after careful LR tuning), opposite to the conclusions of prior work Chen et al. (2024).
RQ2: influence of various model design choices.
Using the multilingual base model further improves fluency and generation in the correct language, but not factuality. Using multilingual IT data only improves the correct language score. Scores for the correct language and fluency get slightly improved for the multilingually pretrained Tower-7B/Dolly-En/FT compared to the similarly-sized English-centric LLaMA-2-7B/Dolly-En/FT. Using multilingual IT data in LLaMA-2-13B/Dolly-DT/FT and Tower-7B/Dolly-DT/FT improves scores for correct language, compared to similar configurations with Dolly-En, but does not improve fluency. Factuality does not get improved with any of the model modifications.
Even though training on small instruction data was shown to be sufficient for English (Zhou et al., 2023), it substantially reduces the cross-lingual capabilities of the final model compared to training on the larger data. The model tuned on (English) LIMA, LLaMA-2-13B/LIMA-En/FT, is characterized by very low scores for all criteria, in non-English evaluation [^1] The helpfulness score assigned by GPT-3.5, 1.52, is substantially higher than the one assigned in human evaluation, 0.94, because LIMA-based model produces much longer outputs than Dolly-based model and GPT-3.5 is known to be biased towards long verbose responses.. This is caused by severe overfitting to English, pronounced by low language correctness scores and generation of incoherent texts. At the same time, scores for English are close to other models, which aligns with the initial findings of (Zhou et al., 2023).
Ablations (small base LLM, LoRA adaptation) reduce scores in non-English. Using LoRA instead of full finetuning, LLaMA-2-13B/Dolly-En/LoRA, and decreasing model size, LLaMA-2-7B/Dolly-En/FT, reduce most of the scores compared to the anchor model LLaMA-2-13B/Dolly-En/FT.
Per-language analysis: fluency is lower for Russian than for French and Portuguese. Per-language analysis presented in App. Figure 4 demonstrates that conclusions discussed above are consistent between languages. A standing-out criteria is fluency which is lower for Russian than for other languages. This is pronounced by the occasional generation of made-up words in Russian and could be connected to the non-Latin script.
@@ 5.1 Main evaluation
Per-task analysis: helpfulness in non-English reduces in some language-related tasks, tasks involving calculation or US-centric factual knowledge. Figure 3 (right) breaks down human-evaluated helpfulness of the anchor model by task category. We find that English-centric model struggles in other languages with some of language-based tasks such as rewriting given sentences, suggesting words that rhyme with the given one or following a given pattern. At the same time, models do succeed on easier language-related tasks such as generate synonyms or words beginning with a given letter. Models also make calculation errors more often in non-English than in English. The low helpfulness for the "sport game" category is connected to the low factuality in non-English: this category asks to explain rules of games popular in the USA and they are explained well in English and often hallucinated in other languages.
5.2 Additional experiment with task modifiers
To complement analysis for RQ1, Tab. 1 reports results on controlling task complexity with task modifiers.
English-centric models are capable of following composite instructions in non-English languages in  65% of cases. The majority of task modifiers are fulfilled in around 80% of cases, with helpfulness score being similar to the value observed in the main evaluation. Interestingly, the instruction to generate response in another language, is fulfilled substantially more often when it is written in non English. An example of the instruction that often fails in non-English is to format the answer as an html page.
Task modifier Mod. fulfill. Helpfulness
en ru en ru
Answer briefly in just a few sentences. 80% 90% 1.70 1.60
Give a detailed answer. 65% 75% 1.60 1.55
List N options (N random from 2 to 10) 66% 83% 1.66 1.66
Answer in X language. (X: Fr, Pt, De) 47% 79% 1.37 1.47
Use markdown formatting in the answer. 92% 100% 1.85 1.28
Format your answer as an html page. 57% 14% 1.35 1.00
Begin each point with the sign –> 7% 14% 0.92 0.85
Capitalize each first letter in the answer. 7% 7% 1.00 0.64
Write in a scientific style. 92% 92% 1.64 1.57
The answer should use simple words. 78% 78% 1.57 1.28
Two-hop instruction, e.g. explain how to serve a dish after telling how to cook it. 93% 86% 1.80 1.60
Average 62% 65% 1.49 1.31
Table 1: Performance with various task modifiers. Modifier fulfilness measures the percentage of inputs for which the modifier was fulfilled. Helpfulness (from 0 to 2) also takes into account the conditions specified in modifiers.
5.3 Preliminary study based on surface metrics
Figure 3 (left) demonstrates surface metrics for all considered model configurations trained with three learning rates, complementing analysis for RQ2.
@@ 5.3 Preliminary study based on surface metrics
Careful hyperparameter tuning and in particular LR selection is essential for achieving multilingual instruction following capabilities. All the model configurations, except training on the small LIMA data, achieve high values for all metrics in all languages with LR of 1e-5 (1e-4 for LoRa adaptation). The lower LR of 1e-6 leads to lower relevance scores in some languages, due to model under-training. On the other side, the higher LR of 1e-4 leads to overfitting to the training language(s), pronounced by lower language correctness scores and lower spellcheck correctness scores, caused by code-switching.
Surface metrics help to select hyperparameters and filter out poor configurations. Surface metrics capture the same effect as in main evaluation, that training on the small LIMA data leads to severe overfitting to English (with all LRs).